OpenStudy (anonymous):

Solve 6z = 1/9 F. z = 54 G. z = 15 H. z = 30 J. z = 1/54

OpenStudy (diyadiya):

Divide both sides by 6, can you?

OpenStudy (anonymous):

6z=1/9 z=1 1 -- * -- 9 6 z=1/54 so option J is the answer

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

you are welcome :)